# RefVNLI: Towards Scalable Evaluation of Subject-driven Text-to-image Generation

### v1

@@ Abstract
Subject-driven text-to-image (T2I) generation aims to produce images that align with a given textual description, while preserving the visual identity from a referenced subject image. Despite its broad downstream applicability—ranging from enhanced personalization in image generation to consistent character representation in video rendering—progress in this field is limited by the lack of reliable automatic evaluation. Existing methods either assess only one aspect of the task (i.e., textual alignment or subject preservation), misalign with human judgments, or rely on costly API-based evaluation. To address this, we introduce RefVNLI, a cost-effective metric that evaluates both textual alignment and subject preservation in a single prediction. Trained on a large-scale dataset derived from video-reasoning benchmarks and image perturbations, RefVNLI outperforms or matches existing baselines across multiple benchmarks and subject categories (e.g., Animal, Object), achieving up to 6.4-point gains in textual alignment and 8.5-point gains in subject consistency. It also excels with lesser-known concepts, aligning with human preferences at over 87% accuracy.
Figure 1: Illustration of RefVNLI: Given a reference image of a subject, a prompt referring to the subject, and a target image, RefVNLI assesses both subject consistency and textual alignment. For subject consistency, it distinguishes identity-preserving variations, like dew on a flower (top image), from identity-altering changes, such as color change (middle image). For textual alignment, it assesses whether the target image reflects all details from the prompt, such as the fence’s position relative to the flower (bottom image).
@@ 1 Introduction
In a well-known scene from “The Little Prince”, the narrator attempts to comfort a grieving prince by saying “I’ll draw you a fence around your flower”. While fairly simple, this offer raises a deeper question: what makes such a drawing adequate? Naturally, it would require precisely following the textual description, i.e., depicting a fence around a flower. However, the use of ‘your’ implies that the drawing must also portray a specific flower—the Prince’s own—one with which he shares a history. Given the flower’s uniqueness and distinctive visual features, the narrator’s task proves far more complex than it first appears.
@@
Figure 2: Qualitative Comparison: We compare RefVNLI with DreamBench++ and CLIP, which score both Subject Consistency (SC) and Textual Alignment (TA), using examples from the Animal, Object, and Human categories. DreamBench++ scores (0-4) are scaled to 0-100 for better readability. RefVNLI demonstrates superior robustness to identity-agnostic changes (SC), such as the zoomed-out parrot (top-middle) and the zoomed-out man with different attire (bottom-middle). It is also more sensitive to identity-defining attributes, assigning low scores to the altered-faced man (bottom-left) and the balloon with mismatched patterns (middle-left). Additionally, RefVNLI excels at detecting text-image mismatches (TA), as seen in its penalization of the top-left image for lacking a waterfall.
@@
Subject-driven text-to-image (T2I) generation [3, 8, 32, 51] enables a variety of downstream applications, such as personalized image generation [51], character consistency in video generation [36], and enhancing vanilla T2I evaluation frameworks for less-known entities via image retrieval. Unlike standard T2I models that are only conditioned on text inputs, subject-driven T2I methods take both a textual prompt and a reference image, enabling more precise subject representation. For example, when creating an image of a fenced-in flower for the Little Prince, such models should use a reference image of the Prince’s flower to ensure the output preserves its unique features (see Fig. 1).
@@
Despite its wide applicability, research in this area has been hindered by the absence of reliable automatic evaluation frameworks. Existing metrics typically correlate poorly with humans, and often focus only on textual alignment between the input prompt and the target image, as in CLIP-T [47] and SigLIP [73], or on subject consistency between the input and target images, as in CLIP-I [47] and DINO-I [7], while both aspects are needed for successful subject-driven generation. More correlative alternatives, like DreamBench++ [46] and VIEScore [27], depend on expensive API calls to models like GPT-4 [42], making them less practical for large-scale research.
To bridge this gap, we present RefVNLI, a cost-effective fine-tuned auto-rater for subject-driven T2I generation. Given a triplet $<$imageref, prompt, imagetgt$>$, RefVNLI predicts two scores—textual alignment and subject consistency—in a single classification, as shown in Fig. 1.
@@
To train RefVNLI, we automatically curate a large-scale dataset of $<$imageref, prompt, imagetgt$>$ triplets, labeled with $<$textual alignment, subject preservation$>$ $\in\{0,1\}^{2}$. For subject preservation, we identify subjects across video frames, creating positive examples using pairs of frames depicting the same subject, and negative ones by pairing frames of different subjects. This approach enables robustness to variations in subject appearance (e.g., rotation, setting, clothing), as well as to the presence of extraneous elements (e.g., dew on the Little Prince’s flower in Fig. 1, top). At the same time, RefVNLI must also be sensitive to identity-defining traits, such as human facial features or object shapes and colors (e.g., middle image in Fig. 1). To this end, we modify images by masking and inpainting identity-critical regions, while keeping everything else unchanged. The original subject crops are then paired with the unaltered images as positive pairs and with the modified images as negative pairs, thereby ensuring the model focuses on key identity attributes.
For textual alignment, we first create positive image-prompt pairs. For that, we use an LLM to caption each image in the aforementioned pairs, ensuring focus on the subject by enclosing it within a bounding box. Negative pairs are then formed by replacing these captions with those of different scenes. To further enhance sensitivity to subtle misalignments, such as a fence drawn next to rather than around the Prince’s flower (Fig. 1, bottom), we also create negative pairs by modifying a single detail in each original (positive) caption, following [12]. Finally, to get the <<<imageref, prompt, imagetgt>>> triplets from each pair of frames and associated captions, we use the cropped subject from one frame as imageref and the entire second frame, alongside its corresponding caption, as {prompt, imagetgt}, resulting in a total of 1.2 million instances.
@@
Figure 3: Generating subject consistency classification training instances from video frames. Given two pairs of frames, each extracted from distinct video scenes featuring the same entity (e.g., a dog), where both frames within each pair depict the same subject (e.g., the same dog), we construct training {imageref, imagetgt} pairs for subject consistency classification. Positive pairs are formed by pairing a cropped subject from one frame (e.g., dog from left frame in Scene 1) with the full frame from the same scene (right frame in Scene 1). In contrast, negative pairs are created by pairing the cropped subject with the other scene’s full frames (e.g., Scene 2). This process is applied to all four frames, with each taking turns as the cropped reference image (imageref), while the corresponding full-frame counterparts serve as imagetgt, yielding a total of 4 positive and 8 negative training pairs.
@@
We evaluate RefVNLI on multiple human-labeled test sets for subject-driven generation, including DreamBench++ [46], ImagenHub [26], and KITTEN [20], across categories such as Humans, Animals, Objects, Landmarks, and a multi-subject setting. For textual alignment, RefVNLI consistently matches or outperforms all baselines, achieving up to 6.4-point gains in Landmarks and excelling at detecting subtle text-image misalignments (e.g., missing waterfall in Fig. 2, top-left). It also shows strong performance for subject preservation, with gains of 6.3 points in the Object category (DreamBench++) and 8.5 points in the multi-subject setting (ImagenHub), surpassing the much larger GPT-4o-based DreamBench++ metric. As seen in Fig. 2, it balances robustness to non-critical changes (e.g., zoomed-out toucan, top-middle) with sensitivity to identity shifts (e.g., altered facial features, bottom-left). Further, as shown in Fig. 6, RefVNLI effectively handles rare subjects (e.g., scientific animal names), aligning with human preferences over 87% of the time.
2 Training Dataset Construction
@@
This section outlines the automated construction process of our training dataset for RefVNLI, comprising triplets of <<<imageref, prompt, imagetgt>>> along with two binary labels: one for the subject preservation of imageref in imagetgt, and another for the textual alignment between the prompt and imagetgt. To that end, we first create subject-driven {imageref, imagetgt} pairs (§2.1) and then auto-generate subject-focused prompts for each imagetgt2.2).
2.1 Subject-driven Image Pairs
To ensure our {imageref, imagetgt} dataset remains robust to identity-agnostic variations (e.g., pose, clothing, lighting, and background changes), we leverage video-based datasets that inherently capture these differences. Specifically, we use Mementos [63], which provides scene-specific video frames with human-written textual descriptions, and TVQA+ [31], which contains human-annotated bounding boxes for characters and objects in TV episodes. We first identify subjects within frames: for Mementos, we extract entities from the provided textual descriptions using Gemini [58] and localize them in the corresponding frames with an object detection model [41], while for TVQA+, we directly use the provided bounding boxes. Positive pairs are formed from frames featuring the same subject, typically within the same scene, [^1] For TVQA+, we also include cross-scene positive pairs for named entities, such as TV characters. while negative pairs consist of frames with distinct subjects, often across scenes, as illustrated in Fig. 3. These frame-pairs are then converted into {imageref, imagetgt}-pairs by cropping the subject from one frame as imageref (e.g., left frame in Scene 1 of Fig. 3) and using the full second frame as imagetgt (each of the other 3 frames in Fig. 3). This is then repeated with reversed roles to create an additional {imageref, imagetgt} pair. In total, we collect 338,551 image pairs (228,661 from Mementos and 109,890 from TVQA+) from 44,418 unique frames.
@@ 2.1 Subject-driven Image Pairs
Figure 4: Creating identity-sensitive {imageref, imagetgt} pairs. Starting with an image and a mask of a subject (e.g., a briefcase), we randomly keep 5 patches within the masked area ([1]) and use them to create 5 inpainted versions ([2]). The version with the highest MSE between the altered and original areas (e.g., bottom image, MSE = 3983) is paired with the unmodified crop to form a negative pair, while the original image and the same crop create a positive pair, with the crop acting as imageref in both cases.
Figure 5: Example of prompt-imagetgt pairs. Given an image with a specific subject (e.g., a dog), we generate a positive prompt by placing a bounding box around the subject and instructing Gemini to describe it (top prompts). Negative prompts are created by swapping positive prompts between images of the same entity (middle prompts). For additional hard negatives, we instruct Gemini to modify a single non-subject detail in the positive prompt while keeping the rest unchanged (bottom prompts).
@@
To further enhance sensitivity to identity-specific attributes, such as facial features in humans or shapes and patterns in objects, we leverage the Open Images dataset [30] to create additional training instances, as shown in Fig. 4. Using its gold segmentation masks, we selectively mask and inpaint identity-critical regions while preserving other details. Specifically, we randomly sample 5 sub-masks covering 30%-50% of the subject mask ([1] in Fig. 4), which we use to create 5 inpainted variants ([2]). The version with the highest Mean Squared Error (MSE) between the modified and original regions (e.g., Fig. 4, bottom image, MSE=3983) is then paired with the unmodified cropped subject to form a negative pair of {imageref, imagetgt}, while the original image and the same crop form a positive pair, with the crop serving as imageref in both cases. This process yields extra 16,572 pairs, helping the model focus on fine-grained identity details. To further improve data quality, we also apply multiple filtering steps, including removing blurry images and those with unclear subjects (see Sec. 9.1 for details).
2.2 Image-prompt Pairs
@@
For each {imageref, imagetgt} pair, we generate positive and negative prompts for imagetgt, as shown in Fig. 5. Positive prompts (Fig. 5, top) are created by instructing Gemini [58] to describe imagetgt, ensuring the subject is explicitly mentioned by enclosing it in a bounding box and guiding the model to focus on it, as well as filtering out prompts where it is missing. For negative prompts (Fig. 5, middle), we swap prompts between frames containing the same type of entity (e.g., a dog). To further enhance sensitivity to subtle mismatches, we also create hard-negative prompts (Fig. 5, bottom) by using Gemini to modify a single non-subject detail in the positive prompts, following [12]. Altogether, combining this step with the image-pairing step (§2.1) yields 1.2 million <<<imageref, prompt, imagetgt>>> triplets labeled with textual alignment and subject preservation.
3 Experimental Settings
This section outlines the training process of RefVNLI3.1), followed by a description of our meta-evaluation protocol and subject-driven benchmarks (§3.2). Lastly, we provide an overview of the baseline models used for comparison (§3.3).
@@ 3.2 Meta-evaluation and Benchmarks
We include 3 subject-driven generation benchmarks with human annotations for textual alignment and subject preservation across various subject categories (e.g., Human, Animal, Object, Landmark). To enable a unified evaluation framework, despite differing scoring methods (5-scale and binary), we convert all annotations into binary labels: one for whether imagetgt fully captures the prompt (textual alignment) and another for whether it correctly depicts the referenced subject (subject preservation). For meta-evaluation, we report the Area Under the ROC Curve (ROC AUC) for each criterion, following standard meta-evaluation practices [18, 70, 72], and also provide a unified score by calculating the harmonic mean of the two ROC AUC scores.
We next detail the 3 datasets included in our analyses.
@@
[46] is a subject-driven generation benchmark with human annotations for 8,190 images generated by 7 models. Annotators rated textual alignment and subject preservation on a 0-4 scale, with each image evaluated by two raters. To convert these ratings into binary labels, we classify a criterion as positive if both annotators assigned a score of at least 3, with at least one giving a 4. We report performance separately for the benchmark’s three subject categories: Human, Animal, and Object. [^3] A fourth ‘style’ category is excluded as it is beyond our work’s scope.
@@
[26] is a standardized evaluation library for conditional image generation models, featuring human annotations across various T2I tasks, including subject-driven generation (150 instances) and multi-concept image composition (102 instances), which involves two referenced subjects per instance. Each image was rated by 3 annotators. Instead of separate ratings for textual alignment and subject preservation, annotators provided a single adherence score (0, 0.5, or 1) per image. To align with our binary labeling framework, images rated 1 by all 3 annotators were assigned positive labels for both criteria, while the rest were re-annotated by this paper’s authors. In the Multi-subject setting, a positive subject preservation label was assigned only when both subjects were accurately depicted. For evaluation, we report performance separately for Animal and Object subjects in the single-subject task. In the Multi-subject task, each instance is split into two single-subject evaluations, with the final score being the lower of the two ratings (for each criterion), to ensure a stricter assessment.
@@
[20] evaluates the ability of T2I models to generate diverse real-world entities, including plants, airplanes, cars, and landmarks. Given 5 reference images and a prompt, annotators rated how well the generated image captured the entity on a 1–5 scale and provided a binary textual alignment score, with each image assessed by five annotators. Unlike our focus on specific subjects, KITTEN evaluates general entity alignment (e.g., a generic rose rather than a specific one). Hence, we only use the 256 Landmark images, as landmarks are unique entities where entity adherence coincides with subject adherence. To convert ratings into binary labels, we apply majority voting for textual alignment and consider subject preservation positive only if most annotators rated it at least 4 and the average score is 4 or higher.

### v2

@@ Abstract
Subject-driven text-to-image (T2I) generation aims to produce images that align with a given textual description, while preserving the visual identity from a referenced subject image. Despite its broad downstream applicability—ranging from enhanced personalization in image generation to consistent character representation in video rendering—progress in this field is limited by the lack of reliable automatic evaluation. Existing methods either assess only one aspect of the task (i.e., textual alignment or subject preservation), misalign with human judgments, or rely on costly API-based evaluation. To address this gap, we introduce RefVNLI, a cost-effective metric that evaluates both textual alignment and subject preservation in a single run. Trained on a large-scale dataset derived from video-reasoning benchmarks and image perturbations, RefVNLI outperforms or statistically matches existing baselines across multiple benchmarks and subject categories (e.g., Animal, Object), achieving up to 6.4-point gains in textual alignment and 5.9-point gains in subject preservation.
[…]
RefVNLI: Towards Scalable Evaluation of
Subject-driven Text-to-image Generation
Aviv Slobodkin1thanks:   Work done during an internship at Google Research.  Hagai Taitelbaum1  Yonatan Bitton1 Brian Gordon1∗Michal Sokolik1Nitzan Bitton Guetta2Almog Gueta1 Royi Rassin1∗Dani Lischinski1Idan Szpektor1 1Google Research   2Ben Gurion University {slobodkin,hagait,yonatanbitton}@google.com
Figure 1: Illustration of RefVNLI: Given a reference image of a subject, a prompt referring to the subject, and a target image, RefVNLI assesses both subject preservation and textual alignment. For subject preservation, it distinguishes identity-preserving variations, like dew on a flower (top image), from identity-altering changes, such as color change (middle image). For textual alignment, it assesses whether the target image reflects all details from the prompt, such as the fence’s position relative to the flower (bottom image).
@@ 1 Introduction
In a well-known scene from “The Little Prince”, the narrator attempts to comfort a grieving prince by saying “I’ll draw you a fence around your flower”. While fairly simple, this offer raises a deeper question: what makes such a drawing adequate? Beyond accurately depicting a fence around a flower, the use of ‘your’ implies that it must portray a specific flower—the Prince’s own— one with which he shares a history. Given the flower’s uniqueness and distinct visual traits, the narrator’s task proves far more complex than it first appears.
@@
Figure 2: Qualitative Comparison: We compare RefVNLI with DreamBench++ and CLIP, which score both Subject Preservation (SP) and Textual Alignment (TA), using examples from the Animal, Object, and Human categories. DreamBench++ scores (0-4) are scaled to 0-100 for better readability. RefVNLI exhibits better robustness to identity-agnostic changes (SP), such as the zoomed-out parrot (top-middle) and the zoomed-out person with different attire (bottom-middle). It is also more sensitive to identity-defining traits, penalizing changed facial features (left-most person) and mismatched object patterns (left and middle balloons). Additionally, RefVNLI excels at detecting text-image mismatches (TA), as seen in its penalization of the top-left image for lacking a waterfall.
@@
Subject-driven text-to-image (T2I) generation Chen et al. (2024); Li et al. (2024); Ruiz et al. (2023) enables a variety of downstream applications, such as personalized image generation Ruiz et al. (2023), character consistency in video generation Liu et al. (2024b), and enhancing vanilla T2I evaluation frameworks for less-known entities via image retrieval (Tahmasebi et al., 2025). Unlike standard T2I models that are only conditioned on text inputs, this setup takes both a textual prompt and a reference image, enabling more precise subject representation. For example, when creating an image of a fenced-in flower for the Little Prince, subject-driven models should use a reference image of the Prince’s flower to ensure the output preserves its unique features (see Fig. 1).
@@
Despite its wide applicability, research in this area has been hindered by a lack of reliable and scalable auto-raters. Existing metrics typically correlate poorly with humans, and often focus only on textual alignment between the input prompt and the target image, as in CLIP-T Radford et al. (2021) and SigLIP Zhai et al. (2023), or on subject preservation between the input and target images, as in CLIP-I Radford et al. (2021) and DINO-I Caron et al. (2021), while both aspects are needed for successful subject-driven generation. More correlative metrics, like DreamBench++ Peng et al. (2024) and VIEScore Ku et al. (2024a), depend on costly API calls to models like GPT-4 OpenAI (2024), making them less scalable and reproducible.
@@
To bridge this gap, we present RefVNLI, a cost-effective fine-tuned auto-rater for subject-driven T2I generation. Given a triplet $<$imageref, prompt, imagetgt$>$, RefVNLI predicts two scores—textual alignment and subject preservation—in a single run, as shown in Fig. 1. To train RefVNLI, we automatically curate a large-scale dataset of $<$imageref, prompt, imagetgt$>$ triplets, labeled with $<$textual alignment, subject preservation$>$ $\in\{0,1\}^{2}$. For subject preservation, we identify subjects across video frames, creating positive examples using pairs of frames depicting the same subject, and negative ones by pairing frames of different subjects (yet of the same entity). This approach enables robustness to variations in subject appearance (e.g., rotation, setting, clothing), as well as to the presence of extraneous elements (e.g., dew on the Little Prince’s flower in Fig. 1, top). At the same time, RefVNLI must also be sensitive to identity-defining traits, such as human facial features or object shapes and colors (e.g., middle image in Fig. 1). To this end, we modify images by masking and inpainting identity-critical regions, while keeping everything else unchanged. The original subject crops are then paired with the unaltered images as positive pairs and with the modified images as negative pairs, thereby teaching the model to focus on key identity attributes.
@@
Figure 3: Generating subject preservation classification training instances from video frames. Given two pairs of frames, each extracted from distinct video scenes featuring the same entity (e.g., a dog), where both frames within each pair depict the same subject (e.g., the same dog), we construct training {imageref, imagetgt} pairs for subject preservation classification. Positive pairs are formed by pairing a cropped subject from one frame (e.g., dog from left frame in Scene 1) with the full frame from the same scene (right frame in Scene 1). In contrast, negative pairs are created by pairing the cropped subject with the other scene’s full frames (e.g., Scene 2). This process is applied to all four frames, with each taking turns as the cropped reference image (imageref), while the corresponding full-frame counterparts serve as imagetgt, yielding a total of 4 positive and 8 negative training pairs.
For textual alignment, we first create positive image-prompt pairs. For that, we use an LLM to caption each image in the aforementioned pairs, ensuring focus on the subject by enclosing it within a bounding box. Negative pairs are then formed by replacing these captions with those of different scenes. For extra sensitivity to minor mismatches, like a fence drawn next to rather than around the Prince’s flower (Fig. 1, bottom), we also create negative pairs by altering a single fact in each original (positive) caption. Finally, to derive the <<<imageref, prompt, imagetgt>>> triplets from each pair of frames and associated captions, we use the cropped subject from one frame as imageref and the entire second frame, alongside its caption, as {prompt, imagetgt}, resulting in a total of 1.2 million instances.
@@
We evaluate RefVNLI on multiple human-labeled test sets for subject-driven generation, including DreamBench++ Peng et al. (2024), ImagenHub Ku et al. (2024b), and KITTEN Huang et al. (2024), across categories such as Humans, Animals, Objects, Landmarks, and a multi-subject setting. For textual alignment, RefVNLI consistently matches or outperforms all baselines, with up to 6.4-point gains in Landmarks and proficiency at detecting subtle text-image misalignments (e.g., missing waterfall in Fig. 2, top-left). It also leads in subject preservation, with gains of up to 6.3 points on Objects (Table 1) and 5.9 points in the multi-subject setting, surpassing the larger GPT-4o-based DreamBench++ baseline. As seen in Fig. 2, it balances robustness to non-critical changes (e.g., zoomed-out toucan, top-middle) with sensitivity to identity shifts (e.g., altered facial features, bottom-left). Further, RefVNLI effectively handles rare subjects (§5), outperforming all baselines and highlighting its value as a reliable alternative to standard T2I metrics for uncommon entities.
2 RefVNLI: Automatic Metric for Subject-driven T2I Generation
We introduce RefVNLI, a cost-effective auto-rater specifically tailored for subject-driven T2I generation. This section details the automated pipeline used to construct its training dataset (§2.1) and the subsequent training process of RefVNLI2.2).
2.1 Training Dataset Construction
@@
To train RefVNLI, we collect a large scale dataset of <<<imageref, prompt, imagetgt>>> triplets, each with two binary labels: one for subject preservation of imageref in imagetgt, and one for textual alignment between the prompt and imagetgt. This involves first creating subject-driven {imageref, imagetgt} pairs, followed by automatic generation of subject-focused prompts for each imagetgt.
@@
Figure 4: Creating identity-sensitive {imageref, imagetgt} pairs. Starting with an image and a mask of a subject (e.g., a bag), we randomly keep 5 patches within the masked area ([1]) and use them to create 5 inpainted versions ([2]). The version with the highest MSE between the altered and original areas (e.g., bottom image, MSE = 3983) is paired with the unmodified crop to form a negative pair, while the original image and the same crop create a positive pair, with the crop acting as imageref in both cases.
Subject-driven image pairs.
To ensure our {imageref, imagetgt} dataset is robust to identity-agnostic changes (e.g., pose, clothing, or lighting changes), we use video-based datasets that inherently capture these differences. Specifically, we use Mementos Wang et al. (2024), comprising scene-specific video frames with human-written textual descriptions, and TVQA+ Lei et al. (2020), containing human-annotated bounding boxes for characters and objects in TV episodes. We first locate subjects within frames: for Mementos, we extract entities from the provided textual descriptions using Gemini Team (2024) and localize them in the associated frames with an object detection model Minderer et al. (2022), while for TVQA+, we directly use the provided bounding boxes. Positive pairs are formed from frames featuring the same subject, usually within the same scene, [^1] For TVQA+, we also include cross-scene positive pairs for named entities, such as TV characters. while negative pairs consist of frames with distinct subjects (of the same type of entity), often across scenes (see Fig. 3). These frame-pairs are then converted into {imageref, imagetgt}-pairs by cropping the subject from one frame as imageref (e.g., left frame in Fig. 3, Scene 1) and using the full second frame as imagetgt (each of the other frames in Fig. 3). This is then repeated with reversed roles for an extra {imageref, imagetgt} pair. In total, we collected 338,551 image pairs (228,661 from Mementos and 109,890 from TVQA+) from 44,418 unique frames.
@@
To further enhance sensitivity to identity-specific attributes, such as facial features in humans or shapes and patterns in objects, we leverage the Open Images dataset Kuznetsova et al. (2020) to create additional training instances, as shown in Fig. 4. Using its gold segmentation masks, we selectively mask and inpaint identity-critical regions while preserving other details. Specifically, we randomly sample 5 sub-masks covering 30%-50% of the subject mask ([1] in Fig. 4), which we use to create 5 inpainted variants ([2]). The version with the highest Mean Squared Error (MSE) between the modified and original regions (e.g., Fig. 4, bottom image, MSE=3983) is then paired with the unmodified cropped subject to form a negative pair of {imageref, imagetgt}, while the original image and the same crop form a positive pair, with the crop serving as imageref in both cases. This process yields extra 16,572 pairs, helping the model focus on fine-grained identity details. To further improve data quality, we also apply multiple filtering steps, including removing blurry images and those with unclear subjects (see Appendix C.2 for details).
Image-prompt pairs.
@@
For each {imageref, imagetgt} pair, we generate positive and negative prompts for imagetgt (Fig. 5). Positive prompts (Fig. 5, top) are created by instructing Gemini Team (2024) to describe imagetgt, ensuring the subject is explicitly mentioned by enclosing it in a bounding box and guiding the model to focus on it, as well as filtering out prompts lacking it. For negative prompts (Fig. 5, middle), we swap prompts between frames containing the same entity type (e.g., a dog). To further enhance sensitivity to subtle mismatches, we also create hard-negative prompts (Fig. 5, bottom) by using Gemini to modify a single non-subject detail in the positive prompts, following Gordon et al. (2024). In total, this and the image-pairing steps yield 1.2 million <<<imageref, prompt, imagetgt>>> triplets labeled for textual alignment and subject preservation.
@@ 2.2 RefVNLI Training
Figure 5: Example of prompt-imagetgt pairs. Given an image with some subject (e.g., a dog), we create a positive prompt by adding a bounding box around the subject and directing Gemini to describe it (top prompts). Negative prompts are created by swapping prompts between images of the same entity (middle prompts). For additional hard negatives, we guide Gemini to modify a single non-subject detail in the positive prompt while keeping the rest unchanged (bottom prompts).
3 Experimental Settings
This section outlines our meta-evaluation protocol and benchmarks (§3.1), followed by an overview of the baseline models used for comparison (§3.2).
@@ 3.1 Meta-evaluation and Benchmarks
We include 3 subject-driven generation benchmarks with human annotations for textual alignment and subject preservation across categories such as Human, Animal, Object, and Landmark. To enable a unified evaluation framework, given differing scoring methods (5-scale and binary), we convert all annotations into binary labels: one for whether imagetgt fully captures the prompt (textual alignment) and another for whether it correctly depicts the referenced subject (subject preservation).
For meta-evaluation, we report ROC AUC for each criterion, following standard practice Honovich et al. (2022); Yarom et al. (2023); Zha et al. (2023), and also compute a unified score as the harmonic mean of the two scores. Following Honovich et al. (2022), significance testing is assessed via bootstrap resampling (Efron, 1987), comparing each baseline to RefVNLI. We report mean scores and highlight models with statistically significant under- or outperformance relative to RefVNLI.
We next present the 3 analyzed benchmarks.
@@
Peng et al. (2024) is a subject-driven generation benchmark with human annotations for 8,190 images generated by 7 models. Annotators rated textual alignment and subject preservation on a 0-4 scale, with each image evaluated by 2 raters. To convert these ratings into binary labels, we classify a criterion as positive if both scores are at least 3, and at least one is a 4. We report performance separately for the benchmark’s three subject categories: Human, Animal, and Object. [^4] A fourth ‘style’ category is excluded as it is beyond our work’s scope.
Textual Alignment Subject Preservation Unified Evaluation
Animal Human Object Animal Human Object Animal Human Object
CLIP   72.8   77.4   74.6   72.4 87.7   76.4   72.6 82.2   75.5
DINO - - - 80.1   78.0   77.3 - - -
Crop-IR - - -   76.9 85.6 83.4 - - -
ArcFace - - - -   61.0 - - - -
CLIPScore   71.5   76.1   72.9 - - - - - -
BLIPScore   75.4   79.5   78.9 - - - - - -
SigLIP   72.5 80.2   77.1 - - - - - -
TIFA   70.6   75.7   69.5 - - - - - -
VQAScore 79.4   78.0 82.6 - - - - - -
VIEScore 77.9   75.2   73.3   63.4 81.1   76.4   69.8 77.7   74.8
DreamBench++ 79.5 82.7 82.5   74.5 84.1   79.4   76.9 83.4   80.9
[…]
RefVNLI 80.2 82.5 82.0 79.4 86.0 85.7 79.8 84.2 83.8
Table 1: ROC AUC scores on DreamBench++ for textual alignment, subject preservation, and their harmonic mean (as a unified evaluation) across Animal, Human, and Object categories. The last two rows feature models finetuned on our dataset, with PaliGemmatext/ref comprising two separate models (PaliGemmatext and PaliGemmaref) trained exclusively for each criterion. Bold indicates the highest score per column. and indicate statistically significant underperformance and outperformance relative to RefVNLI, respectively.
@@
(Ku et al., 2024b) is a human-annotated benchmark for conditional image generation, covering a subject-driven task (150 instances) and a multi-concept task (102 instances), which involves 2 referenced subjects per instance. Each image was rated by 3 annotators. Instead of separate ratings for textual alignment and subject preservation, annotators provided a single adherence score (0, 0.5, or 1) per image. To align with our binary labeling framework, images rated 1 by all 3 annotators were assigned positive labels for both criteria, while the rest were re-annotated by this paper’s authors. In the Multi-subject setting, a positive subject preservation label was assigned only when both subjects were accurately depicted. For evaluation, we report separate scores for Animals and Objects in the single-subject task, while Multi-subject instances are split into two single-subject evaluations, with the final score being the lower rating (per criterion), to ensure a stricter assessment.
@@
Huang et al. (2024) evaluates subject-driven T2I models on generating diverse real-world entities (e.g., plants, vehicles, landmarks), using 5 reference images and a prompt. Annotators rated entity depiction on a 1–5 scale and provided binary textual alignment scores, with each image assessed by 5 annotators. Unlike our focus on specific subjects, KITTEN evaluates general entity alignment (e.g., a generic rose rather than a specific one). Hence, we only use the 256 Landmark images, as landmarks are unique entities where entity adherence coincides with subject adherence. To convert ratings into binary labels, we apply majority voting for textual alignment and consider subject preservation positive only if most annotators rated it at least 4 and the average score is 4 or higher.